Grade 11Mechanics

A motor rotating at 500 rpm is switched off. It comes to rest in 10 seconds. Find the angular retardation and the number of revolutions made by it before stopping.

Adil Majeed
11 Years agoGrade 11

1 Answer

Shobhit Varshney
11 Years ago
Hi,

Given , wo= 500 rpm = 500 /60 rotations per second = 25/3 rps
time = 10 seconds

So, w = 0 = wo+\alphat,
\alpha= – 25/ 30 rotations / s2= -5/6 rotations / s2

Now,\Theta= wo t+ 0.5\alphat2

So,\Theta= (25/3) 10 + 0.5 (-5/6) 102= 125/3 rotations.

thanks